2x-3y+z=2
x-5y+5z=3
3x+y=3z=5
This is a matrices problem
I am having difficutl solving it. I need help.
Thank You
Leticia

Answer by stanbon(75887)

2x-3y+z=2
x-5y+5z=3
3x+y-3z=5
----------------
This is a matrices problem
---
Using the matrix function on my TI-84 I get:
x = 3
y = 2
z = 2
================
Cheers,
Stan H.